For the following scores, find the mean median sum of squared deviations variance and standard deviation. 1382. 1376 1405 1216. 1415
Answer #1

SOLUTION:-

Given data

1382 1376 1405 1216 1415

Here, we have to calculate the  mean :-

Mean=(1382+1376+1405+1216+1415)/5

=6794/5

=1358.8

Mean = 1358.8

Here, we have to calculate the median:-

1216,1376,1382,1405,1415

Median = 1382

Here, we have to calculate the sum of squares of deviations:-

\sum(X-Xbar)^2=26518.8

Here, we have to calculate the variance:-

variance= [(\sumX-Xbar)^2]df

df=4

variance = 26518.8/4

=6629.7

Here, we have to calculate the standard deviation:-

Standard deviation = \sqrtvariance

=81.72296